OpenStudy (aryana_maria2323):

What is the best substitution to make to evaluate the integral of the quotient of cosine of 2 times x and the square root of the quantity 5 minus 2 times the sine of 2 times x, dx? a. u=sin(2x) b. u=cos(2x) c. u=2x d. u=5-2sin(2x)

OpenStudy (xguardians):

Can you type it out (sorry I always request this).

OpenStudy (aryana_maria2323):

No your fine. I always forget to do that. One second

Nnesha (nnesha):

\[\int\limits_{}^{}\frac{ \cos(2x) }{ \sqrt{5-2\sin(2x)}} dx \] like this ?

OpenStudy (aryana_maria2323):

Yes exactly lol

OpenStudy (xguardians):

I'm always bad with these types of substitutions, I'd say work with the denominator and take out d. u=5-2sin(2x) Let me double check.

Nnesha (nnesha):

well.. the point of u-substitution to cancel out something and makes the messy stuff looks easier if i pick top part for u and then take the derivative u=cos(2x) du= -2sin (2x) \[\int\limits_{}^{}\frac{ \cos(2x) }{ \sqrt{5-2\sin(2x)}} dx \]  that doesn't make any sense. I can't cancel out anything so u should be whatever under the sqrt